OpenStudy (anonymous):

Let f(x) = 4x – 32 and g(x) = 4x. Part 1 What is (f ◦ g)(x) ? Part 2 Describe the method you used to solve this problem. I want to double check if my answers are right, but I don't understand part 2?

OpenStudy (lalaly):

(f ◦ g)(x) = f(g(x)) f(4x)=4(4x)-32 = 16x-32 part2: g(x) is the x ....that means you use g(x) instead of x i guess

OpenStudy (anonymous):

So for the first part you are like distributing the x to f and g?

OpenStudy (lalaly):

(f ◦ g)(x) is equivalent to f ( g(x) ) u substitute the whole g(x) instead of x in f(x) since g(x)=4x so f( g(x) )= f(4x) = 4(4x) - 32

OpenStudy (anonymous):

Hehe, don't believe me eye ;)

OpenStudy (anonymous):

Lol I do, I just want a further explanation xD

OpenStudy (anonymous):

Ok now I understand, thank you guys so much :)